Case: In December 2004, R. E. Torgler was trying to decide whether to add a new line of injection
molded plastic products to those already manufactured and distributed by Reto S.A. In order to do
so, the firm would have to buy new injection molding equipment; none of the existing equipment
could be adapted to perform the necessary operations, and Torgler was anxious to retain control of
manufacturing. Actually, new injection molding equipment of the type needed had been considered
before, but a decision to purchase the equipment had been postponed because the product concept
was judged to need additional development. But now the products seemed ready.

Sales of the new product were forecast at SFr. 2,000,0001 per year, from which a sales commission
of 15 percent would be paid to Reto’s sales agents. Direct manufacturing costs were budgeted at SFr.
600,000 for materials and SFr. 900,000 for labor, leaving an annual cash flow before taxation of SFr.
200,000. The new equipment would cost SFr. 600,000, delivered and installed, and was expected to
have an economic life of 10 years before it would become worthless.

Reto was able to borrow money at 8 percent, although it did not plan to negotiate a loan
specifically for the purchase of this equipment.

Questions:

1. Ignoring the effect of taxes, what is the “internal rate of return” (IRR) on the proposed
investment. Assume the new equipment would be installed by January 1, 2005, and begin
producing on that date.
2. The cost of the equipment can be deducted from annual cash flows before they are subjected
to taxation. Assuming that the equipment will last 10 years, and that an equal amount of the
cost of SFr. 600,000 will be deducted each year, and that the tax rate is expected to be 45
percent, what is the IRR on an after-tax basis?
3. Torgler has stated that Reto should be willing to purchase this machine as long as it yielded a
return of 12 percent after taxation. Should he make the investment?
4. Actually, to stimulate industrial development, the tax rules allow up to one third of the cost of
any such investment to be deducted from cash flows in the first year after the investment, and
up to one-fifth of the remainder of the unamortized investment amount can be deducted in
the second year. Thereafter, annual deductions are computed on a straight-line basis such that
no more than the original cost of the equipment is depreciated over its economic life. How, if
at all, does this tax rule affect the attractiveness of the investment?
Answer #1
Sales Forecast            20,00,000
Sales Commissions              3,00,000 15%
Direct manufacturing costs
Materials              6,00,000
Labour              9,00,000
Annual cashflow before Tax              2,00,000

1)

Annual cashflow before Tax              2,00,000
P.V.A.F @ 8% for 10 Years                     6.71
P V of Cash Inflows            13,42,000
PV of Cash outflows              6,00,000
NPV @8%              7,42,000 NPV1
Annual cashflow before Tax              2,00,000
P.V.A.F @ 35% for 10 Years                     2.72
P V of Cash Inflows              5,44,000
PV of Cash outflows              6,00,000
NPV @35%                -56,000 NPV2
IRR= R1+(NPV1*(R2-R1))/(NPV1-NPV2)
IRR= 8+ (742000*(35-8))/(742000+56000)
IRR= 8+ 25.10526316
IRR= 33.105%

2)

Annual cashflow before Tax              2,00,000
Less: Depreciation                 60,000
Profit before Taxes              1,40,000
Less: Taxes                 63,000 45%
Profit after Taxes                 77,000
Add: Depreciation                 60,000
Annual cashflow after Tax              1,37,000
P.V.A.F @ 8% for 10 Years                     6.71
P V of Cash Inflows              9,19,270
PV of Cash outflows              6,00,000
NPV @8%              3,19,270 NPV1
Annual cashflow after Tax              1,37,000
P.V.A.F @ 35% for 10 Years                     2.72
P V of Cash Inflows              3,72,640
PV of Cash outflows              6,00,000
NPV @35%            -2,27,360 NPV2
IRR= R1+(NPV1*(R2-R1))/(NPV1-NPV2)
IRR= 8+ (319270*(35-8))/(319270+227360)
IRR= 8+ 15.76988091
IRR= 23.770%

3)

Profit after Taxes=                 77,000
Cost of Machine=              6,00,000
Rate of return after taxes= 12.83%
Yes he should make the investment as he is getting 12.83% return on the machine

4)

1st Year 2nd Year 3-10 Years
Annual cashflow before Tax              2,00,000                              2,00,000        2,00,000
Less: Depreciation              2,00,000                                  80,000           40,000
(1/3 of 600,000) 1/5 of (600,000-200,000) (1/8*(600,000-200,000-80,000)
Profit before Taxes                          -                                1,20,000        1,60,000
Less: Taxes                          -                                    54,000           72,000 45%
Profit after Taxes                          -                                    66,000           88,000
Add: Depreciation              2,00,000                                  80,000           40,000
Annual cashflow after Tax              2,00,000                              1,46,000        1,28,000
P.V.A.F @ 8%                     0.93                                      0.86 4.92
(6.71-0.93-0.86)
P.V of Inflows        1,85,185.19                         1,25,171.47 6,29,760.00
P.V of Inflows              9,40,117
PV of Cash outflows              6,00,000
NPV @8%              3,40,117 NPV1
1st Year 2nd Year 3-10 Years
Annual cashflow after Tax              2,00,000                              1,46,000        1,28,000
P.V.A.F @ 35%                     0.74                                      0.55 1.43
(2.72-0.74-0.55)
P V of Cash Inflows              1,48,148                                  80,110        1,83,040
P V of Cash Inflows              4,11,298
P V of Cash Outflows              6,00,000
NPV @35%            -1,88,702 NPV2
IRR= R1+(NPV1*(R2-R1))/(NPV1-NPV2)
IRR= 8+ (340117*(35-8))/(340117+188702)
IRR= 8+ 17.36541047
IRR= 25.365%

This Tax rule makes investment more beneficial as more depreciation is allowed in first two years where time value of money is more as compared to later years and so IRR increases to 25.365% as compared to 23.77%

Please like the solution if satisfied with the answer and if any issues please mention it in comments...thanks